Phosphorus pentachloride decomposes according to the chemical equation

PCl5(g)−⇀↽−PCl3(g)+Cl2(g) Kc=1.80 at 250

A 0.1414 mol sample of PCl5(g) is injected into an empty 2.00 L reaction vessel held at 250 ∘C.

Calculate the concentrations of PCl5(g) and PCl3(g) at equilibrium.
Answer #1

Solution:-

Initial [PCl5] = moles / volume = (0.1414 mol) / (2.00 L) = 0.0707 M

PCl5 (g) PCl3 (g) Cl2 (g)
Intial 0.0707   0   0
Change -x   ​​+x +x
Equilbrium (0.0707 - x)   x x

Now,

Kc = [PCl3] [Cl2] / [PCl5]

1.80 = (x.x) / (0.0707 - x)

1.80 = x2 / (0.0707 - x)

0.1273 - 1.80x = x2

​​​​​​or,

x2 + 1.80x - 0.1273 = 0 ................................(i)

Using quadratic formula,

±x = [(-b) ± √(b2 - 4ac)] / 2a

From eq.(i) ,

a = 1 ; b = 1.80 ; c = -0.1273

±x = [(-1.80) ± √(1.80)2 - 4 (1) (-0.1273)] / 2 (1)

±x = [(-1.80) ± √(3.34 + 0.509)] / 2  

±x = [(-1.80) ± 1.93623] / 2

x = +0.0681 or, x = -1.868

Using positive value of x,

x = 0.0681 M

Thus, at equilibrium,

[PCl3 (g)] = x = 0.0681 M

[PCl5 (g)] = (0.0707 - x) = (0.0707 - 0.0681) = 0.0026 M

---------------------------------------------------------------